OpenStudy (anonymous):

anyone familiar with wolfram alpha? how do i get it show/solve the equation a=v/t for v=15 and t=4?

OpenStudy (anonymous):

a=15/4 just divide

OpenStudy (anonymous):

i know but my assignment is to do one complete question showing all steps using wolfram alpha so i need it show the formula and also the units

sam (.sam.):

type 'find a, a=v/t , v=15 , t=4'

OpenStudy (anonymous):

i still don't get it to show the answer with that..

OpenStudy (anonymous):

u want more steps?? i dont think there's much more to it.

OpenStudy (anonymous):

i just need to get to show 3.75...but it only gives 15/4

OpenStudy (anonymous):